a si solar cell 2cm ×2cm width Is=10^-9 has an optical. generation rate of 10^18 EHP/cm^3 with Ln=Lp=2 micrometer.if the depletion width is 1 micrometer . calculate the short circuit current and open circuit voltage for cell
1
Expert's answer
2019-11-18T10:38:33-0500

First, find the photocurrent, which is equal to the short circuit current:

"I_{op} = q\cdot A\cdot g_{op}\cdot (L_p+L_n+W) , where A=2\cdot 2 cm^2"

"I_{op} = 1.6\cdot10^{-19}\cdot 4\cdot10^{-4}\cdot 10^{24}\cdot 10^{-6}\cdot (2+2+0.5)=0.288mA"

Then, by the known short circuit current, we find the open circuit voltage:

"V_{oc} = \frac{kT}{q}\cdot ln(\frac{I_{op}}{I_{th}} + 1)"

"V_{oc} = 0.025\cdot ln(\frac{0.288\cdot10^{-3}}{10^{-9}} + 1) = 0.487 V"